Q3: Mike decides to take the mortgage loan to buy a house with total price of $200,000.He made 50,000 for down payment. He decided to pay back the money every quarter in the equal amount. What should be his equal quarterly payment be over the next 20 years if the annual interest rate is 7%?
Answer #1

Total Loan Amount = 200,000 -50,000 = 150,000
Rate per quarter = 7%/4
Number of Periods = 20*4 = 80

PMT per quarter = PV/(1-(1+r)-n)/r = 150,000/(1-(1+7%/4)-80)/(7%/4) = 3498.14

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well